OpenStudy (spyro517):

Point G is located at (3, −1), and point H is located at (−2, 3). Find the x value for the point that is 2 over 3 the distance from point G to point H. 1.5 −0.33 −0.5 6.33

OpenStudy (spyro517):

adam.bloedel

OpenStudy (spyro517):

@adam.bloedel

OpenStudy (anonymous):

I got B

OpenStudy (spyro517):

alright thank you for the help.

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

@herua123 If you're going to answer, please explain in detail. It will help him more than just giving an answer. Have logic behind your answer.

OpenStudy (anonymous):

@Spyro517 Do you understand how he arrived at B? Or would you like me to go through how to find the answer again?

OpenStudy (spyro517):

@adam.bloedel Can you go into detail of how he arrived at his answer please?

OpenStudy (anonymous):

Differece in x form G to H 3 - -2 = 5 2/3 of the distance 2/3 * 5 = 10/3 From H to 2/3 of the way to H 3 - 10/3 = -1/3

OpenStudy (spyro517):

Well that actually explains it perfectly. Thank you for proving your statement.

OpenStudy (anonymous):

+1 @herua123

OpenStudy (anonymous):

:) and your welcome:))